OpenStudy (anonymous):

Is (4, 6) a solution to y = x – 2? Explain.

hero (hero):

y = x - 2 6 = 4 - 2 6 = 2 False

OpenStudy (anonymous):

if you replace the (4, 6) coordinates into the equation x = x - 2, do you get a true statement?

OpenStudy (anonymous):

plug in you x and y to the problem. y = 6 and x = 4 2- 4 = 6

OpenStudy (anonymous):

*** y = x - 2